Given the function f(x)= |x+4|, find each of the following
f(7), f(-8), f(0)

Answer by fractalier(6550)

Merely plug in these values for x, and get
f(7) = |7 + 4| = 11
f(-8) = |-8 + 4| = |-4| = 4
f(0) = |0 + 4| = 4